### v1

Abstract
In this paper, we address the issue of estimating and inferring the distributional treatment effects in randomized experiments. The distributional treatment effect provides a more comprehensive understanding of treatment effects by characterizing heterogeneous effects across individual units, as opposed to relying solely on the average treatment effect. To enhance the precision of distributional treatment effect estimation, we propose a regression adjustment method that utilizes the distributional regression and pre-treatment information. Our method is designed to be free from restrictive distributional assumptions. We establish theoretical efficiency gains and develop a practical, statistically sound inferential framework. Through extensive simulation studies and empirical applications, we illustrate the substantial advantages of our method, equipping researchers with a powerful tool for capturing the full spectrum of treatment effects in experimental research.
Keywords: Randomized experiment, A/B testing, Distributional treatment effect, Heterogeneous treatment effect, Regression adjustment, Distributional regression
JEL Codes: C14, C32, C53, E17, E44
1 Introduction
Since the foundational works of Fisher (1925, 1935) and Neyman (1923), randomized controlled trials (RCTs) have become essential in evaluating the effectiveness of treatments or interventions. Over the last few decades, RCTs, also known as A/B testing, have emerged as key tools for causal inference across a wide range of scientific and engineering disciplines. These fields include medicine (Rubin, 1997), economics (Duflo et al., 2007; Athey and Imbens, 2017), political science (Imai, 2005; Horiuchi et al., 2007), and sociology (Baldassarri and Abascal, 2017), among others. For a more thorough exploration, refer to the modern textbook treatments in Imbens and Rubin (2015).
In RCTs, researchers often encounter the challenge of low detection power, where the variance of the treatment effect estimator is large relative to the actual treatment effect, leading to ambiguous results (e.g., Lewis and Rao, 2015). Contrary to the common belief that large sample sizes alone can resolve this issue, research indicates this is not always the case (Deng et al., 2013). Increasing sample size or experiment duration is costly and not always feasible. Therefore, using estimators with smaller variance can improve precision and lead to more reliable estimates. Additionally, examining distributional treatment effects, not just average treatment effects, is of practical interest to researchers (Bitler et al., 2006).
In this paper, we introduce a regression adjustment method to improve the precision of estimating the distributional treatment effect (DTE) in RCTs. The proposed method uses pre-treatment information through distributional regression (DR). Our approach addresses the challenge of low detection power, employing a semiparametric approach in that the underlying distributions are left unspecified. This method accommodates a wide range of outcome distributions, including discrete, continuous, and mixed variables, and is useful for documenting heterogeneous treatment effects.
We theoretically demonstrate that regression adjustment can reduce the variance of distribution estimators and enhance the precision of the DTE estimator. We establish the uniform asymptotic properties of the DTE estimator. We also introduce a practical resampling method for statistical inference and establish its theoretical validity. Through Monte Carlo simulations, we evaluate our method and show that regression adjustment significantly improves the precision of the DTE estimator and outperforms existing methods. Additionally, we apply our method in two empirical studies: a randomized experiment on the impact of nudges on water usage and the Oregon Health Insurance Experiment. Our results highlight the importance of distributional treatment estimation with regression adjustment.
Our paper provides several important contributions to the existing literature. First, our paper contributes to a growing body of literature on heterogeneous treatment effects in econometrics and statistics (e.g., Athey and Imbens, 2016; Wager and Athey, 2018). The quantile treatment effect (QTE) was first introduced by Doksum (1974) and Lehmann (1975) for estimating treatment effects due to unobserved heterogeneity. Since then, various estimation and inference methods for the distributional and quantile treatment effect have been developed and applied, including Heckman et al. (1997), Abadie et al. (2002), Athey and Imbens (2006), Bitler et al. (2006), Chernozhukov and Hansen (2005), Djebbari and Smith (2008), Donald and Hsu (2014), Callaway et al. (2018), Callaway and Li (2019), Chernozhukov et al. (2020), and Kallus et al. (2024), among others. Jiang et al. (2023) studies the estimation of regression-adjusted QTE under covariate adaptive randomization schemes. To the best of our knowledge, however, techniques for estimating DTE has not been explored in the context of regression-adjustment method.666 To estimate heterogeneous treatment effects, researchers commonly use two mututally not exclusive approaches: conditional average treatment effect (CATE) and distributional treatment effect (DTE). CATE estimates the average treatment effect for subgroups based on observed variables (Athey and Imbens, 2016; Imai and Ratkovic, 2013; Künzel et al., 2019; Nie and Wager, 2020; Wager and Athey, 2018). DTE, on the other hand, compares outcome distributions between treatment and control groups across the entire spectrum, and thus it can capture effects that CATE might miss, particularly those stemming from unobserved factors. As such, DTE remains valuable even after accounting for observed variables in CATE analyses.
Next, the literature has extensively investigated the use of pretreatment covariates to estimate ATE with the regression adjustment (e.g. Fisher, 1925; Cochran, 1957; Cox and McCullagh, 1982; Frison and Pocock, 1992). Freedman (2008b, a) and Lin (2013) have noted the potential issue of the adjustment due to finite-sample bias and misspecification. Recent works by Yang and Tsiatis (2001), Tsiatis et al. (2008) and Ansel et al. (2018) apply linear regression adjustment, while Rosenblum and Van Der Laan (2010) and Bartlett (2018) explore nonlinear regression adjustment. List et al. (2022) use machine learning methods for regression adjustment. Imbens and Rubin (2015) and Negi and Wooldridge (2021, 2020) consider regression adjustment for both linear and nonlinear regression models. These existing works focus on average treatment effect (ATE), whereas our research broadens the scope of regression adjustment to estimate the impact of treatment on the entire distribution in randomized experiments.
Thirdly, for our regression adjustment, we utilize the DR approach, which can be considered a semiparametric approach for estimating conditional distributions, in that both marginal and conditional distributions are left unspecified. The semiparametric property is essential in practice since we rarely know the true underlying distributions in many applications. Williams and Grizzle (1972) introduce DR to analyze ordered categorical outcomes by using multiple binary regressions. Foresi and Peracchi (1995) extend the approach to characterize a continuous, conditional distribution and Hall et al. (1999) propose a local version of distributional regression. Chernozhukov et al. (2013) establish its uniform validity for estimating the entire conditional distribution when an asymptotical continuum of binary regressions is used. See also Chernozhukov et al. (2020), Chernozhukov et al. (2023), Wang et al. (2023, 2024) for later developments.
The rest of the paper is organized as follows. Section 2 introduces the potential outcomes framework and defines the treatment effect parameters of interest within our study. Section 3 presents theoretical results. In Section 4, we provide finite-sample properties through Monte Carlo simulations. Section 5 offers two empirical case studies. Finally, we conclude in Section 6. All proofs and additional simulation results can be found in Supplemental Material.
2 Treatment Effect and Estimation Method
This section first introduces the setup and explain a model for regression adjustment. In what follows, we use $\mathbb{E}[\cdot]$ and $\mathrm{Var}(\cdot)$ as the expectation and variance operator, respectively. All small-$o$ notation and asymptotic results, including $o_{p}(1)$ and $O_{p}(1)$, in the paper are relative to sample size $n$. Let $\ell^{\infty}(T)$ be the collection of all bounded real-valued functions defined on an arbitrary index set $T$. We use $\rightsquigarrow$ to denote convergence in law. That is, given $X_{n}\in\ell^{\infty}(T)$ or $X_{n}(t)$ for $t\in T$, $X_{n}\rightsquigarrow X$ in $\ell^{\infty}(T)$ if $X_{n}$ converge in law to $X$ in $\ell^{\infty}(T)$.
2.1 Setup and Treatment Effects
Consider a randomized controlled trial with $K$ treatments, where each individual is randomly assigned to one of the treatments with pre-specified probabilities $\{\pi_{k}\}_{k=1}^{K}$ with $\sum_{k=1}^{K}\pi_{k}=1$ and $\pi_{k}>0$ for all $k\in\mathcal{K}:=\{1,\dots,K\}$. We consider the potential outcome framework (Neyman, 1923; Rubin, 1974). Let $Y$ denote the (observed) outcome of interest with its support $\mathcal{Y}\subset\mathbb{R}$ and $Y(k)$ the potential outcome under treatment $k\in\mathcal{K}$. Then the following relation holds: $Y=\sum_{k=1}^{K}W_{k}Y(k)$, where $W_{k}$ is set to 1 if an individual is assigned to treatment $k\in\mathcal{K}$ and 0 otherwise and satisfies $\Pr(W_{k}=1)=\pi_{k}$ and $\sum_{k=1}^{K}W_{k}=1$.
Suppose we observe a random sample $\{(\bm{W}_{i},\bm{X}_{i},Y_{i})\}_{i=1}^{n}$ drawn independently from the distribution of $(\bm{W},\bm{X},Y)$ with a sample size of $n$. Here, $\bm{W}:=(W_{1},\dots,W_{K})^{\top}$ denotes the vector of treatment assignment indicators for $K$ treatments with its support $\mathcal{W}:=\{0,1\}^{K}$ and $\bm{X}$ a $p\times 1$ vector of pre-treatment covariates with its support $\mathcal{X}\subset\mathbb{R}^{p}$. We assume that $\bm{X}$ realizes before the randomized controlled trials and is independently and identically distributed even across treatment statuses. Given $\bm{W}_{i}=(W_{1,i},\dots,W_{K,i})^{\top}$, the number of observations assigned to treatment $k\in\mathcal{K}$ is defined as $n_{k}:=\sum_{i=1}^{n}W_{k,i}$ and the empirical probability of treatment assignment is given by $\hat{\pi}_{k}:=n_{k}/n$ for $k=1,\dots,K$.
One prevalent approach to measure the treatment effect is the average treatment effect (ATE), defined as $ATE_{k,k^{\prime}}:=\mathbb{E}[Y(k)]-\mathbb{E}[Y(k^{\prime})]$ for $k,k^{\prime}\in\mathcal{K}$. ATE is simple to estimate and interpret, while it is incapable of estimating the treatment effect on the entire distribution. To consider the distributional features, let $F_{Y(k)}(y)$ be the distribution function of $Y(k)$ for treatment status $k\in\mathcal{K}$ and $\gamma:=(\mathbb{F}_{Y(1)},\dots,\mathbb{F}_{Y(K)})^{\top}\in\Gamma\subseteq\ell^{\infty}(\mathcal{Y})^{K}$ a vector of the potential outcome distributions. We define the distributional treatment effect (DTE) between treatments $k,k^{\prime}\in\mathcal{K}$ as
$\Delta^{DTE}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{DTE}(\gamma),$
where the functional $\Psi_{k,k^{\prime}}^{DTE}:\Gamma\to\ell^{\infty}(\mathcal{Y})$ is defined as $\Psi_{k,k^{\prime}}^{DTE}(\gamma)=F_{Y(k)}-F_{Y(k^{\prime})}$. One of its advantages is that DTE is well-defined for any type of outcome, including discrete variables.777 Our approach measures the distributional effect as the difference between the distributions of potential outcomes under distinct treatments. While this paper focuses on DTE, we note that under the rank invariance assumption, our results are also informative about the distribution of individual treatment effects (ITE). Without additional conditions or rank invariance, researchers may consider adopting a partial-identification approach for estimating the distribution of ITE. See Heckman et al. (1997) among others.
Moreover, the distributional information is an essential building block for other types of treatment effect parameters. For example, fixing some constant $h>0$, we can write the probability $\Pr\{y<Y(k)\leq y+h\}=F_{Y(k)}(y+h)-F_{Y(k)}(y)$, thereby defining the probability treatment effect (PTE) as
$\Delta^{PTE}_{k,k^{\prime},h}:=\Psi_{k,k^{\prime},h}^{PTE}(\gamma),$
where the functional $\Psi_{k,k^{\prime},h}^{PTE}:\Gamma\to\ell^{\infty}(\mathcal{Y})$ is defined as
$\Psi_{k,k^{\prime},h}^{PTE}(\gamma)(y)$ $=$ $(F_{Y(k)}-F_{Y(k^{\prime})})(y+h)-(F_{Y(k)}-F_{Y(k^{\prime})})(y).$
PTE measures changes in the probability that the outcome variable falls in interval $(y,y+h]$ and can apply for any type of outcome variables. For an integer outcomes and $h=1$, $\Delta_{k,k^{\prime}}^{PTE}(y,1)$ quantifies the probability change at location $y$, or $\Pr\{Y(k)=y\}-\Pr\{Y(k^{\prime})=y\}$. Alternatively, when the outcome variable is continuous, we can consider QTE, defined through a simple inversion of the distribution function: for $u\in(0,1)$, $\Delta^{QTE}_{k,k^{\prime}}(u):=F_{Y(k)}^{-1}(u)-F_{Y(k^{\prime})}^{-1}(u),$ where $u\mapsto F_{Y(k)}^{-1}(u)$ is the quantile function of $Y(k)$. Since standard analysis based on quantile functions presumes the existence of the outcome density, QTE is not readily applicable in the presence of discreteness in the outcome distributions.
2.2 Distributional Regression
In this paper, we extend the idea of a regression-adjustment approach for estimating the DTE and PTE, using the distributional regression (DR) framework with pre-treatment covariates. The true conditional distribution of $Y(k)$ given $\bm{X}$ can be written as
$F_{Y(k)|\bm{X}}(y|\bm{X})=\mathbb{E}[1\!{\rm l}_{\{Y(k)\leq y\}}|\bm{X}],$
[…]
In many practical applications, the true conditional distribution $F_{Y(k)|\bm{X}}(y|\bm{X})$ remains unknown. It is crucial not to impose overly restrictive global parametric constraints. To address this issue, we employ the distribution regression framework, wherein a parametric linear-index model is used to describe the binary outcomes $1\!{\rm l}_{\{Y(k)\leq y\}}$ for each $y\in\mathcal{Y}$. More specifically, letting $\Lambda:\mathbb{R}\to[0,1]$ be a known link function, we define the following conditional distribution based on a linear index model for each fixed $y\in\mathcal{Y}$:
$G_{Y(k)|\bm{X}}(y|\bm{X}):=\Lambda\big{(}T(\bm{X})^{\top}\beta_{k}(y)\big{)},$ (1)
where $T:\mathcal{X}\to\mathbb{R}^{d}$ is some transformation, such as polynomial or pair-wise interaction of regressors, specified by researchers, and $\beta_{k}(y){\in}\mathbb{R}^{d}$ are unknown parameters. Setting the link function as the normal or logistic distribution function, we can consider the models as probit or logit models for binary outcome $1\!{\rm l}_{\{Y(k)\leq y\}}$, respectively.
It is worth noting that the unknown parameters $\beta_{k}(y)$ are specific to point $y\in\mathcal{Y}$, and can be considered as pseudo-parameters to characterize the true conditional distribution at each point $y$. These pseudo-parameters capture local information of the distributions of interest, and can be estimated at the parametric rate under certain regularity conditions, as detailed later. Thus, we view the conditional distribution model $G_{Y(k)|\bm{X}}(\cdot|\bm{X})$ in (1) as an approximation model for the true conditional distribution $F_{Y(k)|\bm{X}}(\cdot|\bm{X})$, and our subsequent analysis accommodates potential model misspecifications.
We estimate the DR model in (1) using the quasi-maximum likelihood framework (see White, 1982), separately for observations exposed to different treatments. For each $(k,y)\in\mathcal{K}\times\mathcal{Y}$, the estimator for $\beta_{k}(y)$ is defined as the maximizer of the log-likelihood function:
$\widehat{\beta}_{k}(y)=\arg\max_{\beta\in\mathcal{B}}\widehat{\ell}_{k}(\beta;y),$ (2)
where ^$\widehat{\ell}_{k}(\beta;y):=n_{k}^{-1}\sum_{i=1}^{n}W_{k,i}\cdot\ell_{i}(\beta;y)$ is the log-likelihood function for treatment $k\in\mathcal{K}$, and $\mathcal{B}\subset\mathbb{R}^{d}$ is the parameter space. Given the estimator $\widehat{\beta}_{k}(y)$, we define the conditional distribution function estimator as
$\widehat{G}_{Y(k)|\bm{X}}(y|\bm{x}):=\Lambda\big{(}T(\bm{x})^{\top}\widehat{\beta}_{k}(y)\big{)}.$ (3)
We can estimate the conditional distribution above for each value of $y\in\mathcal{Y}$ when the outcome variable $Y$ takes finite discrete values, and we select a sufficiently large set of discrete points $\{y_{j}\in\mathcal{Y}\}_{j=1}^{J}$ for estimation when the outcome variable is a continuous random variable.888 One important property of the conditional distribution function is non-decreasing by definition, while the estimator $\widehat{G}_{Y(k)|\bm{X}}(y|\bm{x})$ does not necessarily satisfy monotonicity in finite samples due to estimation errors. We can monotonize the conditional distribution estimators using the rearrangement method, for example, as proposed by Chernozhukov et al. (2009).
2.3 Regression-Adjusted Treatment Effect Estimator
By randomization, the potential outcome distributions $F_{Y(k)}$ for $k\in\mathcal{K}$ are identified as the outcome distributions under treatment $k$. Hence, it can be estimated from the data $\{(\bm{W}_{i},\bm{X}_{i},Y_{i})\}_{i=1}^{n}$. The simple DTE and PTE estimators are based on the empirical distribution functions $\widehat{\mathbb{F}}_{Y(k)}^{simple}(y):=n_{k}^{-1}\sum_{i=1}^{n}W_{k,i}\cdot 1\!{\rm l}_{\{Y_{i}\leq y\}}$ for $k\in\mathcal{K}$. We propose a new estimator based on the distributional regression adjustment with pre-treatment covariates. Define the discrete probability measures ^$\widehat{\mathbb{P}}_{\bm{X}}:=n^{-1}\sum_{i=1}^{n}\delta_{\bm{X}_{i}}$ and ^$\widehat{\mathbb{P}}_{\bm{X}}^{(k)}:=n_{k}^{-1}\sum_{i=1}^{n}W_{k,i}\cdot\delta_{\bm{X}_{i}}$ for covariates $\bm{X}$ and for $k\in\mathcal{K}$, where $\delta_{\bm{x}}$ is the measure that assigns mass 1 at $\bm{x}\in\mathcal{X}$. For an arbitrary real-valued function $f:\mathcal{X}\to\mathbb{R}$, we denote by
^$\widehat{\mathbb{P}}_{\bm{X}}f=\frac{1}{n}\sum_{i=1}^{n}f(\bm{X}_{i})\ \ \ \ \mathrm{and}\ \ \ \ \widehat{\mathbb{P}}_{\bm{X}}^{(k)}f=\frac{1}{n_{w}}\sum_{i=1}^{n}W_{k,i}\cdot f(\bm{X}_{i}).$
Then, for treatment $k\in\mathcal{K}$, we define the regression-adjusted distribution function as,
$\widehat{\mathbb{F}}_{Y(k)}:=\widehat{\mathbb{P}}_{\bm{X}}\widehat{G}_{Y(k)|\bm{X}}\ \ \ \mathrm{in}\ \ell^{\infty}(\mathcal{Y}).$ (4)
The regression-adjusted estimator is the sample average of conditional distribution functions over pre-treatment covariates utilizing all observations, where these conditional distributions are estimated using only observations pertaining to specific treatments.
Letting $\widehat{\gamma}:=(\widehat{\mathbb{F}}_{Y(1)},\dots,\widehat{\mathbb{F}}_{Y(K)})^{\top}$, we define the regression-adjusted DTE and PTE estimators:
$\widehat{\Delta}^{DTE}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{DTE}(\,\widehat{\gamma}\,)\ \ \ \mathrm{and}\ \ \ \widehat{\Delta}^{PTE}_{k,k^{\prime},h}:=\Psi_{k,k^{\prime},h}^{PTE}(\,\widehat{\gamma}\,).$ (5)
The linear index model in (3) can be considered as one of the generalized linear models (GLM). In the literature, it is well known that the certain choice of specifications, such as a link function, can guarantee the unbiased estimator for the unconditional mean in finite samples. Such choice of the link function is called the canonical link function (see McCullagh and Nelder, 1989). More precisely, the canonical link function should yield
$\frac{1}{n_{k}}\sum_{i=1}^{n}W_{k,i}\cdot\big{(}1\!{\rm l}_{\{Y_{i}\leq y\}}-\widehat{G}_{Y(k)|\bm{X}}(y|\bm{X}_{i})\big{)}=0.$ (6)
At the population level, this relation can be expressed as: $\mathbb{E}\big{[}1\!{\rm l}_{\{Y(k)\leq y\}}-G_{Y(k)|\bm{X}}(y|\bm{X})\big{]}=0.$ This property is important for the regression adjustment to guarantee that the regression adjustment does not induce bias to improve the precision of the treatment effect estimator. Negi and Wooldridge (2021) pointed this out for a nonlinear regression adjustment method and derive the asymptotic property when the true distribution is known. In what follows, we focus on the canonical class of the GLM model for binary outcome, which includes OLS and logit, while we do not assume any global parametric assumptions and allow for model mis-specification.
3 Theoretical Results
In this section, we first derive the oracle efficiency of the proposed estimator. We subsequently present asymptotic properties and a resampling method.
3.1 Oracle Estimator and Efficiency
In this section, we investigate possible efficiency gain from the regression adjustment, when the population counterpart related to the adjustment is available. This analysis will allow us to establish the efficiency property.
Leveraging the finite-sample unbiasedness property in (6), we can rewrite (4) as
$\widehat{\mathbb{F}}_{Y(k)}:=\widehat{\mathbb{F}}_{Y(k)}^{simple}+\big{(}\widehat{\mathbb{P}}_{\bm{X}}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\big{)}\widehat{G}_{Y(k)|\bm{X}},$
where the second term of the right-hand side involves the DR estimator for the adjustment. In the above equations, we meanwhile consider the situation where the estimator $\widehat{G}_{Y(k)|\bm{X}}$ is replaced by the population model $G_{Y(k)|\bm{X}}(y|\bm{X})$ in (1). Then, define the oracle distribution function with the regression adjustment
$\widetilde{\mathbb{F}}_{Y(k)}$ $:=\widehat{\mathbb{F}}_{Y(k)}^{simple}+\big{(}\widehat{\mathbb{P}}_{\bm{X}}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\big{)}G_{Y(k)|\bm{X}}.$
Theorem 1 below presents the efficiency gain resulting from the application of our regression adjustment method, given the oracle estimator above. Although this oracle estimator is infeasible, the subsequent section (specifically, in the proof of Theorem 2) establishes an asymptotic equivalence between the feasible and oracle estimators.
For Theorem 1, we make the following assumptions regarding the experimental setup and sampling scheme:
Assumption 1.
(a) The observations $\{(\bm{W}_{i},\bm{X}_{i},Y_{i})\in\{0,1\}^{K}{\times}\mathcal{X}{\times}\mathcal{Y}\}_{i=1}^{n}$ are $n$ independent copies of the random variables $(\bm{W},\bm{X},Y)$.
(b) Treatment assignment is independent of both potential outcomes and pre-treatment covariates: $\bm{W}\perp\!\!\!\perp\big{(}Y(1),\dots,Y(K),\bm{X})$. The pre-specified treatment probabilities $\{\pi_{k}\}_{k=1}^{K}$ satisfy $\pi_{k}>0$ for every $k\in\mathcal{K}$ and $\sum_{k=1}^{K}\pi_{k}=1$.
Assumption 1 (a) requires random sampling and Assumption 1 (b) is standard under a randomized controlled trial with multiple treatments.
Theorem 1.
(a) Suppose that Assumption 1 holds and that $\hat{\pi}_{k}=\pi_{k}+o(1)$ as $n\to\infty$ for every $k\in\mathcal{K}$. Then, for each $k\in\mathcal{K}$ and for any $y\in\mathcal{Y}$, we have in $\ell^{\infty}(\mathcal{Y})$,
$\mathrm{Var}\big{(}\hat{\mathbb{F}}_{Y(k)}^{simple}\big{)}\geq\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}+o(n^{-1}),$
where the equality holds only if $G_{Y(k)|\bm{X}}=F_{Y(k)}$.
(b) Define $\widehat{\gamma}^{simple}:=(\widehat{\mathbb{F}}_{Y(1)}^{simple},\dots,\widehat{\mathbb{F}}_{Y(K)}^{simple})^{\top}$ and $\widetilde{\gamma}:=(\widetilde{\mathbb{F}}_{Y(1)},\dots,\widetilde{\mathbb{F}}_{Y(K)})^{\top}$. Additionally, we assume that the model specification for distributional regression in (1) is correct. Then,
$\mathrm{Var}\big{(}\widehat{\gamma}^{simple}\big{)}\succeq\mathrm{Var}\big{(}\widetilde{\gamma}\big{)}+o(n^{-1}),$
where $\succeq$ denotes the positive semi-definiteness. When $\mathrm{Var}\big{(}F_{Y(k)}(y|\bm{X})-r\cdot F_{Y(k^{\prime})}(y|\bm{X})\big{)}>0$ for any distinct $k,k^{\prime}\in\mathcal{K}$ and $r\in\mathbb{R}$, the positive definite result holds.
Theorem 1(a) shows the efficiency gains achieved by applying regression adjustment to estimate unconditional distribution functions. More specifically, the variance of the unconditional distribution estimators decreases if $G_{Y(k)|\bm{X}}(y)$ has some predictive power for $\mathbbm{1}_{\{Y(k)\leq y\}}$. Furthermore, Theorem 1(b) elaborates on these gains in a positive semi-definite sense for a vector of regression-adjusted estimators. This result encompasses any linear combination of the unconditional distribution estimators and implies potential efficiency gains for both DTE and PTE estimators. While a similar positive-definiteness result for non-linear regression adjustment is reported in Theorem 6 of Negi and Wooldridge (2020), our result provides an additional insight that the efficiency gains occur when the regression models are not linearly dependent across treatments. Theorem 1(b) requires correct specification, which is analogous to the requirement in Negi and Wooldridge (2020) for non-linear regression adjustment.
3.2 Asymptotic Distribution of DTE Estimator
In this subsection, we derive the asymptotic distribution of the DTE and PTE estimators, which is useful for statistical inference. We also introduce a resampling technique and show its asymptotic validity. We obtain the results in this section under the following assumptions:
Assumption 2.
For any $y\in\mathcal{Y}$ and $k\in\mathcal{K}$, the following conditions hold.
(a) The log-likelihood functions $\beta\mapsto\ell_{k}(\beta;y)$ is a concave function. The link function $\Lambda(\cdot)$ belongs to the canonical class and is twice continuously differentiable with its first derivative $\lambda(\cdot)$.
(b) The true parameters $\beta_{k}(y)$ uniquely solve the maximization problem in $\max_{\beta}\mathbb{E}[\widehat{\ell}_{k}(\beta;y)]$ and are contained in the interior of the compact parameter space $\mathcal{B}$.
(c) The maximum eigenvalues of $H_{k}(y)$ is strictly negative uniformly over $y\in\mathcal{Y}$.
(d) If $Y$ is a continuous random variable, then the conditional density function $f_{Y|\bm{X}}(y|\bm{x})$ exists, is uniformly bounded, and is uniformly continuous in $(y,x)\in\mathcal{Y}\times\mathcal{X}$.
Assumption 2 (a)-(c) are needed for deriving the asymptotic property of the DR estimators and are standard for the likelihood estimation of binary-choice models in the literature. Assumption 2 (d) is required only when the outcome variable is continuous.
In the below theorem, we derive the limiting distribution of the regression-adjusted DTE and PTE estimators.
Theorem 2.
Suppose that Assumptions 1 and 2 hold. Then, we have, in $\ell^{\infty}(\mathcal{Y})$,
$\sqrt{n}(\widehat{\Delta}^{DTE}_{k,k^{\prime}}-\Delta^{DTE}_{k,k^{\prime}})\rightsquigarrow\Psi^{DTE}_{k,k^{\prime}}(\mathbb{B}),$
$\sqrt{n}(\widehat{\Delta}^{PTE}_{k,k^{\prime},h}-\Delta^{PTE}_{k,k^{\prime},h})\rightsquigarrow\Psi^{PTE}_{k,k^{\prime},h}(\mathbb{B}),$
for every $k,k^{\prime}\in\mathcal{K}$ and some constant $h>0$, where $\mathbb{B}$ is a mean-zero Gaussian process defined in Lemma A.4 of Supplemental Material.
The limiting processes in the above theorem rely solely on the sampling error of the original data, $\{(\bm{W}_{i},\bm{X}_{i},Y_{i})\}_{i=1}^{n}$, and are free from estimation errors in DR. This is ensured through the use of the stochastic equicontinuity property in the theorem’s proof, which is in line with the argument presented in Newey (1994). This result highlights the significance of our DR-based regression adjustment, as it does not add additional variance to the DTE and PTE estimators for large sample sizes.
Although it is simplified, the limiting processes of the DTE and PTE estimators depend on unknown nuisance parameters and may complicate inference in finite samples. To address the challenge of non-pivotal limiting processes in finite samples, we turn to the exchangeable bootstrap (as described in Praestgaard and Wellner, 1993; van der Vaart and Wellner, 1996). This approach consistently estimates limit laws of relevant empirical distributions, allowing us to consistently estimate the limit process of the estimator via the functional delta method.
For the resampling scheme, we introduce a vector of random weights $(S_{1},\dots,S_{n})$. To establish the validity of the bootstrap, we assume that the random weights satisfy the following assumption.
Assumption 3 (Bootstrap).
Let $(S_{1},\dots,S_{n})$ be $n$ scalar, nonnegative random variables, which are independent and identically distributed and also independent of the original sample. The random weights $\{S_{i}\}_{i=1}^{n}$ satisfy the followings: $\mathbb{E}|S_{i}|^{2+\epsilon}<\infty$ for some $\epsilon>0$, $\bar{S}_{n}:=n^{-1}\sum_{i=1}^{n}S_{i}\to^{p}0$ and $n^{-1}\sum_{i=1}^{n}(S_{i}-\bar{S}_{n})^{2}\to^{p}1$.
This resampling scheme encompasses a variety of bootstrap methods, such as the empirical bootstrap, subsampling, wild bootstrap and so on (See van der Vaart and Wellner, 1996).
Under treatment $k\in\mathcal{K}$, the empirical distribution of the outcome variable with the random weights and the empirical probability measure of $\bm{X}$ with the random weights are respectively given by
$\widehat{\mathbb{F}}_{Y(k)}^{simple\ast}(y):=\frac{1}{n_{k}}\sum_{i=1}^{n}S_{i}\cdot W_{k,i}\cdot 1\!{\rm l}_{\{Y_{i}\leq y\}}\ \ \ \ \mathrm{and}\ \ \ \ \widehat{\mathbb{P}}_{\bm{X}}^{\ast}:=\frac{1}{n}\sum_{i=1}^{n}S_{i}\cdot\delta_{\bm{X}_{i}}.$
Then, for treatment $k\in\mathcal{K}$, we define regression-adjusted distribution functions in $\ell^{\infty}(\mathcal{Y})$ as,
$\widehat{\mathbb{F}}_{Y(k)}^{\ast}:=\widehat{\mathbb{P}}_{\bm{X}}^{\ast}\widehat{G}_{Y(k)|\bm{X}}^{\ast}.$ (7)
In the above equation, we use $\widehat{G}_{Y(k)|\bm{X}}^{\ast}$ the conditional distribution estimator based on the original sample, instead of the bootstrap sample. According to Theorem 2, the conditional distribution estimator has no impact on the limit distribution and we keep the estimator throughout the resampling procedure. The validity is proven in the theorem below.
Letting $\widehat{\gamma}^{\ast}:=(\widehat{\mathbb{F}}_{Y(1)}^{\ast},\dots,\widehat{\mathbb{F}}_{Y(K)}^{\ast})^{\top}$, we define the bootstrapped treatment effect estimator with regression adjustment as
$\widehat{\Delta}^{DTE\ast}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{DTE}(\,\widehat{\gamma}^{\ast}\,)\ \ \ \mathrm{and}\ \ \ \widehat{\Delta}^{PTE\ast}_{k,k^{\prime},h}:=\Psi_{k,k^{\prime},h}^{PTE}(\,\widehat{\gamma}^{\ast}\,).$
In the theorem provided below, we show that the exchangeable bootstrap provides a method to consistently estimate the limit process of the regression-adjusted DTE and PTE estimators, following van der Vaart and Wellner (1996).
Theorem 3.
Under Assumptions 1-3, we have, in $\ell^{\infty}(\mathcal{Y})$,
[…]
Theorems 2 and 3 present a methodologically rigorous and empirically useful approach for estimating DTE and PTE. Moreover, these theoretical results establish the asymptotic validity of the exchangeable bootstrap procedure for constructing confidence intervals and conducting hypothesis tests for both DTE and PTE.
4 Monte-Carlo Simulation Results
For our Monte Carlo simulation, we consider various setups to cover continuous/discrete and symmetric/skewed outcome distributions. We report our results from four types of data generating processes (DGPs): DGP1-DGP4. In all DGPs, we use the same setup for pre-treatment regressors $\bm{X}=(1,X_{1},X_{2})^{\top}$ with a uniform random variable $X_{1}\sim U(0.5,1.5)$ and a normal random variable $X_{2}\sim N(0,1)$. The treatment status $W_{1}$ follows the binomial distribution with $\Pr\{W_{1}=1\}=\pi_{1}\in\{0.3,0.5\}$.
Under DGP1 and DGP2, we generate continuous outcomes from the following model:
$Y=X_{1}+(X_{1}+X_{2})\cdot W_{1}+|X_{1}+X_{2}|\cdot U,$
where the unobserved component $U$ is specified as $U\sim N(0,1)$ under DGP1 and $U\sim\chi_{3}^{2}$ under DGP2, respectively. Consequently, the unobserved variable is symmetrically distributed under DGP1, while it exhibits a skewed distribution under DGP2. Under DGP3 and DGP4, we consider discrete outcome distributions satisfying the following conditional mean:
$\mathbb{E}[Y|\bm{X},\bm{W}]=\exp(W_{1}+X_{1}+X_{2}/2).$
Given the conditional mean model above, we set the outcome distribution conditional on $\bm{X}$ and $\bm{W}$ to follow the Poisson distribution under DGP3 and to follow the negative binomial distribution with the dispersion parameter $r=5$ under DGP4, respectively.
Figure 1: Performance metrics of simple and regression-adjusted DTE estimators
(DGP1, continuous outcome, $\pi_{1}=0.5$)
Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
Figure 2: Performance metrics of simple and regression-adjusted DTE estimators
(DGP3, discrete outcome, $\pi_{1}=0.5$): Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
We consider three estimators: the simple DTE (Simple) and the DTE with regression adjustment based on the linear regression (OLS) and logit model (Logit). To evaluate the finite sample performance of these DTE estimators, we compute their bias and root mean squared error (RMSE) across 10,000 simulations. Additionally, we assess the average length and coverage probability of their 95% pointwise confidence intervals. In the main text, we present results for DGP1 and DGP3 under a treatment assignment probability of $\pi_{1}=0.5$. Figures displaying results for all other combinations of DGPs and treatment assignment probabilities can be found in Supplemental Material B.
Figure 1 summarizes the performance metrics for DGP1, while Figure 2 presents results for DGP3 under a treatment assignment probability of $\pi_{1}=0.5$. For continuous outcomes in DGP1 and DGP2, we consider thresholds $y$ at quantiles $\{0.1,\dots,0.9\}$ of the true outcome distribution. For discrete outcomes in DGP3 and DGP4, we consider $y\in\{1,\dots,5\}$. The results are displayed using boxplots over these threshold locations for a concise visual representation. To approximate true distributions, we generate datasets with 1,000,000 observations and compute the DTE at each threshold $y$.
As illustrated in Figure 1, both the RMSE and the average length of the confidence intervals decrease as the sample size $n$ increases, as expected. The regression-adjusted estimators improve upon the simple estimator by reducing the RMSE and the length of confidence intervals by approximately 13%-25% across all sample sizes. However, for smaller sample sizes ($n\in\{300,500\}$), the regression-adjusted estimators exhibit a slight undercoverage, with a coverage probability around 0.945. Similar trends are observed with the discrete outcome depicted in Figure 2. In this scenario, the regression-adjusted estimators shorten the confidence intervals by about 9%-20% compared to the simple estimator.
In summary, these findings suggest that while regression adjustment does not affect the bias in DTE estimation, it significantly enhances precision. Additionally, we confirm that regression adjustment reduces the length of confidence intervals, thereby enhancing precision, while maintaining their 95% coverage probability.
5 Empirical Applications
5.1 Nudge and Water Consumption
In this section, we reanalyze data from a randomized experiment conducted by Ferraro and Price (2013b) in 2007 to examine the impact of norm-based messages, or nudges, on water usage in Cobb County, Atlanta, Georgia. [^9] The dataset is publicly available as Ferraro and Price (2013a) at https://doi.org/10.7910/DVN1/22633. The experiment implemented and compared three different interventions aimed at reducing water usage against a control group (no nudge). For more details on the experiment and the results of the average treatment effect analysis, refer to Ferraro and Price (2013b). For simplicity, following List et al. (2022), we focus exclusively on the nudge that demonstrated the strongest average effect over the control group, which combined prosocial appeal with social comparison. We estimate the regression-adjusted DTE and PTE of this intervention over the control group, using the monthly water consumption in the year prior as pre-treatment covariates $\bm{X}$. The outcome variable $Y$ represents the level of water consumption from June to September 2007, measured in thousands of gallons and discretely distributed. It is important to note that, while the measure in gallons appears approximately continuous, the presence of subtle discreteness can complicate both theoretical and practical statistical inference. Consequently, QTE is not applicable in this context.
Figure 3: Nudge Effect on Water Consumption (thousands of gallons)
Distributional and Probability Treatment Effect
Note: Top left: simple DTE; top right: regression-adjusted DTE (logit model). Bottom left: simple PTE; bottom right: regression-adjusted PTE (logit model). Shaded areas and error bars: 95% pointwise confidence intervals. $n=78,500$.
Figure 3 presents the DTE and PTE of the intervention in comparison to the control group. We compute the DTE for $y\in\{0,1,2,\dots,200\}$. The top left panel of Figure 3 displays the simple estimate of the DTE, whereas the top right panel illustrates the regression-adjusted estimate of the DTE obtained with logit model. The shaded areas denote the 95% pointwise confidence intervals. The results indicate that the regression-adjusted DTE has around 4% to 32% smaller standard errors and consequently tighter confidence intervals compared to the simple DTE.
The bottom left panel of Figure 3 depicts the simple PTE, while the bottom right panel shows the regression-adjusted PTE, both with 95% pointwise confidence intervals. The PTE is presented in increments of $h=10$ for $y\in\{0,10,20,\dots,200\}$. The findings suggest that the treatment is effective in reducing the probability of high water consumption and increasing the probability of low water consumption. Specifically, the regression-adjusted PTE results indicate an probability increase in water usage in the range of (20,30] and (30, 40] by 1.67 percentage points (pp) with standard errors (s.e.) of 0.42 pp and 1.31 pp (s.e. = 0.46 pp), respectively. On the other hand, it suggests a probability decrease in water usage in the range of (60,70], (70, 80], (90, 100] and (100, 110] by 0.61 pp (s.e. = 0.24 pp), 0.41 pp (s.e. = 0.20 pp), 0.35 pp (s.e. = 0.15 pp), and 0.26 pp (s.e. = 0.12 pp), respectively. The variance reduction is particularly notable in the range between 70 and 110. The probability change for the range (70,80], (90,100], and (100,110] are not significant under simple estimates, but are significantly negative under regression adjustment. [^10] Some significance levels are difficult to discern in Figure 3 due to the scale. The probability of decreased water consumption in the range $(80,90]$ is 0.33 pp with a standard error of 0.17 pp under regression-adjusted PTE, despite appearing significantly negative. Additionally, for the ranges $(90,100]$ and $(100,110]$, the probabilities of decreased water consumption are 0.28 pp (s.e. = 0.15 pp) and 0.22 pp (s.e. = 0.12 pp), respectively, under simple PTE.
5.2 Insurance Coverage and Healthcare Utilization
In this subsection, we investigate the impact of insurance coverage on healthcare utilization using data from the Oregon Health Insurance Experiment. [^11] The dataset is publicly available at https://www.nber.org/research/data/oregon-health-insurance-experiment-data. In 2008, the state of Oregon offered insurance coverage to a group of uninsured low-income adults through a lottery. Twelve months after the lottery, a mail survey was conducted to collect data on various outcomes. The treatment in this experiment was randomly assigned based on the number of people in the household. Therefore, we focus on individuals who reported being in single-person households, excluding those who included one or two additional people in their lottery application. Refer to Finkelstein et al. (2012) for more details about the experiment and the treatment effect results on various outcomes.
We examine healthcare utilization by focusing on the number of primary care (or doctor) visits over a six-month period, denoted as our outcome variable $Y$, collected from survey data. Our set of covariates $\bm{X}$ includes the applicant’s age, gender, race, and income. Not everyone who was selected by the lottery to apply for insurance coverage ended up obtaining it. Previous research by Finkelstein et al. (2012) has documented a positive intention-to-treat (ITT) effect and a local average treatment effect (LATE) of insurance coverage on the number of primary care visits. In this illustrative example, we use the lottery assignment (rather than actual insurance enrollment) as our treatment variable and focus solely on estimating the ITT effect. To obtain the local distributional treatment effect, one can use the representation provided in Abadie (2002), applying the ITT effect at each location $y$.
Figure 4: Effect of Health Insurance Coverage on Doctor Visits
Distributional and Probability Treatment Effect
Note: Top left: simple DTE; top right: regression-adjusted DTE (logit model). Bottom left: simple PTE; bottom right: regression-adjusted PTE (logit model). Shaded areas and error bars: 95% pointwise confidence intervals. n=13,759.
Figure 4 shows the distributional effect of insurance coverage on the number of primary care visits. The DTE estimates (top left) indicate a statistically significant effect on visits ranging from 0 to 6. The regression-adjusted DTE estimates obtained with logit model are nearly identical to the simple DTE estimates and reduce the standard errors by 0.1%-1.1% across values of $y{\in}\{0,1,{\dots},30\}$. The gains from regression adjustment are modest in this case.
In the bottom panel of Figure 4, we present the estimated PTE with 95% confidence intervals for each integer value between 0 and 15 (i.e., $h=1$). The regression-adjusted PTE estimates suggest that insurance coverage reduces the probability of not seeing a doctor by approximately 6.56 pp (s.e. = 0.82 pp) and increases the probability of seeing a doctor three, four, five, and six times by 2.36 pp (s.e. = 0.61 pp), 0.66 pp (s.e. = 0.31 pp), 0.96 pp (s.e. =0.28 pp), and 0.81 pp (s.e. = 0.31 pp), respectively.
6 Conclusion
In this paper, we introduce a regression adjustment method to improve the precision of estimating distributional treatment effects through the use of distributional regression. We analyze the theoretical properties of our proposed method and provide a practical inference method that has been proven to be theoretically valid. Our simulation results demonstrate the remarkable performance of our method. Additionally, we apply our method to real-world data in two empirical applications and uncover heterogeneous treatment effects. In both applications, we find clearer conclusions compared to existing methods.
Supplementary Material
In this supplementary material, we complete the proofs in Section A, and provide additional results from our simulation experiment in Section B. The relevant R code is available online.
Appendix A Theoretical Results
A.1 Proof of Theorem 1
To ensure the completeness of the proof, we begin by presenting a variant of Lagrange’s identity and Bergström’s inequality in the below lemma, which is useful to prove the efficiency gain of the regression adjustment.
Lemma A.1.
For any $(a_{1},\dots,a_{K})\in\mathbb{R}^{K}$ and $(b_{1},\dots,b_{K})\in\mathbb{R}^{K}$ with $b_{k}>0$ for all $k=1,\dots,K$, we can show that
$\sum_{k=1}^{K}\frac{a_{k}^{2}}{b_{k}}-\frac{\big{(}\sum_{k=1}^{K}a_{k}\big{)}^{2}}{\sum_{k=1}^{K}b_{k}}=\frac{1}{\sum_{k=1}^{K}b_{k}}\cdot\frac{1}{2}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq k\end{subarray}}^{K}\frac{(a_{k}b_{\ell}-a_{\ell}b_{k})^{2}}{b_{k}b_{\ell}},$
which implies Bergström’s inequality, given by
$\sum_{k=1}^{K}\frac{a_{k}^{2}}{b_{k}}\geq\frac{\big{(}\sum_{k=1}^{K}a_{k}\big{)}^{2}}{\sum_{k=1}^{K}b_{k}}.$
Proof.
Lagrange’s identity is that, for any $(c_{1},\dots,c_{K})\in\mathbb{R}^{K}$ and $(d_{1},\dots,d_{K})\in\mathbb{R}^{K}$,
$\bigg{(}\sum_{k=1}^{K}c_{k}^{2}\bigg{)}\bigg{(}\sum_{k=1}^{K}d_{k}^{2}\bigg{)}-\bigg{(}\sum_{k=1}^{K}c_{k}d_{k}\bigg{)}^{2}=\frac{1}{2}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq k\end{subarray}}^{K}(c_{k}d_{\ell}-c_{\ell}d_{k})^{2}.$ (A.1)
Fix arbitrary $(a_{1},\dots,a_{K})\in\mathbb{R}^{K}$ and $(b_{1},\dots,b_{K})\in\mathbb{R}^{K}$ with $b_{k}>0$ for all $k=1,\dots,K$. Then, taking $c_{k}=a_{k}/\sqrt{b_{k}}$ and $d_{k}=\sqrt{b_{k}}$ for all $k=1,\dots,K$ in (A.1), we can show that
$\bigg{(}\sum_{k=1}^{K}\frac{a_{k}^{2}}{b_{k}}\bigg{)}\bigg{(}\sum_{k=1}^{K}b_{k}\bigg{)}-\bigg{(}\sum_{k=1}^{K}a_{k}\bigg{)}^{2}$ $=\frac{1}{2}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq k\end{subarray}}^{K}\bigg{(}\frac{a_{k}}{\sqrt{b_{k}}}\sqrt{b_{\ell}}-\frac{a_{\ell}}{\sqrt{b_{\ell}}}\sqrt{b_{k}}\bigg{)}^{2}$
$=\frac{1}{2}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq k\end{subarray}}^{K}\frac{(a_{k}b_{\ell}-a_{\ell}b_{k})^{2}}{b_{k}b_{\ell}},$
which leads to the desired equality. Also, the last expression in the math display above is non-negative, which leads to Bergström’s inequality. ∎
Proof of Theorem 1.
Part (a) Choose any arbitrary $k\in\mathcal{K}$. The quadratic expansion of $\widetilde{\mathbb{F}}_{Y(k)}=\widehat{\mathbb{F}}_{Y(k)}^{simple}-(\widehat{\mathbb{P}}_{X}^{(k)}-\widehat{\mathbb{P}}_{X})G_{Y(k)|\bm{X}}$ yields
$\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}=\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}-2\mathrm{Cov}\Big{(}\widehat{\mathbb{F}}_{Y(1)}^{simple},(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-\widehat{\mathbb{P}}_{\bm{X}})G_{Y(k)|\bm{X}}\Big{)}$
$\hskip 85.35826pt+\mathrm{Var}\Big{(}(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-\widehat{\mathbb{P}}_{\bm{X}})G_{Y(k)|\bm{X}}\Big{)}.$ (A.2)
We can write ^$\widehat{\mathbb{P}}_{\bm{X}}=\sum_{k\in\mathcal{K}}\hat{\pi}_{k}\widehat{\mathbb{P}}_{\bm{X}}^{(k)}$. It is assumed that observations are a random sample and $n_{k}/n=\pi_{k}+o(1)$ for every $k\in\mathcal{K}$ as $n\to\infty$. Furthermore, all unconditional and conditional functions are bounded. By applying the dominated convergence theorem, we can show
$n\mathrm{Cov}\Big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple},(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-\widehat{\mathbb{P}}_{\bm{X}})G_{Y(k)|\bm{X}}\Big{)}$ $=n\mathrm{Cov}\Big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple},(1-\hat{\pi}_{k})\widehat{\mathbb{P}}_{\bm{X}}^{(k)}G_{Y(k)|\bm{X}}\Big{)}$
$=\frac{1-\pi_{k}}{\pi_{k}}\mathrm{Cov}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}},G_{Y(k)|\bm{X}}\big{)}+o(1).$ (A.3)
Similarly, we can show that
$n\mathrm{Var}\big{(}(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-\widehat{\mathbb{P}}_{\bm{X}})G_{Y(k)|\bm{X}}\big{)}$ =$n\mathrm{Var}\big{(}(1-\hat{\pi}_{k})\widehat{\mathbb{P}}_{\bm{X}}^{(k)}G_{Y(k)|\bm{X}}\big{)}+n\sum_{\ell:\ell\not=k}\mathrm{Var}\big{(}\hat{\pi}_{\ell}\widehat{\mathbb{P}}_{\bm{X}}^{(\ell)}G_{Y(k)|\bm{X}}\big{)}$
=$\frac{(1-\pi_{w})^{2}}{\pi_{w}}\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}+\sum_{\ell:\ell\not=k}\frac{\pi_{\ell}^{2}}{\pi_{\ell}}\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}+o(1)$
=$\bigg{(}\frac{(1-\pi_{k})^{2}}{\pi_{k}}+\sum_{\ell:\ell\not=k}\pi_{\ell}\bigg{)}\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}+o(1)$
$=\frac{1-\pi_{k}}{\pi_{k}}\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}+o(1).$ (A.4)
It follows from (A.2)-(A.4) that
$n\big{\{}\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}-\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}\big{\}}$
$\ \ \ \ \ \ =\frac{1-\pi_{k}}{\pi_{k}}\big{\{}2\mathrm{Cov}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}},G_{Y(k)|\bm{X}}\big{)}-\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}\big{\}}+o(1).$ (A.5)
Applying addition and subtraction, we obtain
$2\mathrm{Cov}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}},G_{Y(k)|\bm{X}}\big{)}-\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}$ $=\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}\big{)}-\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)},$
which together with (A.1) yields
$n\big{\{}\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}-\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}\big{\}}$ $=\frac{1-\pi_{k}}{\pi_{k}}\Big{\{}\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}\big{)}-\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}\Big{\}}+o(1).$
Since $G_{Y(k)|\bm{X}}$ belongs to the canonical class, we have that $\mathbb{E}[G_{Y(k)|\bm{X}}]=F_{Y(k)}$ and thus we can show that $\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}=\mathbb{E}\big{[}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}^{2}\big{]}$. Here, the equality hold only when $F_{Y(k)|\bm{X}}=F_{Y(k)}$ or $\bm{X}$ has no predictive power for the event $1\!{\rm l}_{\{Y(k)\leq\cdot\}}$.
Part (b) Under the correct specification, an application of the law of iterated expectation yields that $\mathrm{Cov}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}},G_{Y(k)|\bm{X}}\big{)}=\mathrm{Var}\big{(}\mathbb{E}[1\!{\rm l}_{\{Y(k)\leq\cdot\}}|\bm{X}]\big{)},$ which together with (A.1) shows
$n\big{\{}\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}-\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}\big{\}}$ $=\frac{1-\pi_{k}}{\pi_{k}}\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}+o(1).$
Since $\pi_{k}\in(0,1)$ and $\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}\big{)}\geq 0$, it follows that $\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}\geq\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}+o(n^{-1}).$
Now, we shall show that, for any $k,\ell\in\mathcal{K}$,
$n\mathrm{Cov}\big{(}\widehat{\mathbb{F}}_{Y(k)},\widetilde{\mathbb{F}}_{Y(\ell)}\big{)}=\mathrm{Cov}\big{(}G_{Y(k)|\bm{X}},G_{Y(\ell)|\bm{X}}\big{)}.$ (A.6)
Fix any two distinct treatment statuses $k,\ell\in\mathcal{K}$. We can write $\widetilde{\mathbb{F}}_{Y(k)}=\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}G_{Y(k)|\bm{X}}\big{)}+\widehat{\mathbb{P}}_{\bm{X}}G_{Y(k)|\bm{X}}$ and also ^$\widehat{\mathbb{P}}_{\bm{X}}G_{Y(k)|\bm{X}}=\sum_{\ell\in\mathcal{K}}\hat{\pi}_{\ell}\widehat{\mathbb{P}}_{\bm{X}}^{(\ell)}G_{Y(k)|\bm{X}}$. Given random sample and the bi-linear property of the covariance function, we can show that
$\mathrm{Cov}\Big{(}\widetilde{\mathbb{F}}_{Y(k)},\widetilde{\mathbb{F}}_{Y(\ell)}\Big{)}=$ $\mathrm{Cov}\Big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}G_{Y(k)|\bm{X}},\hat{\pi}_{\ell}\widehat{\mathbb{P}}_{\bm{X}}^{(\ell)}G_{Y(\ell)|\bm{X}}\Big{)}$
$+\mathrm{Cov}\big{(}\hat{\pi}_{k}\widehat{\mathbb{P}}_{\bm{X}}^{(k)}G_{Y(k)|\bm{X}},\widehat{\mathbb{F}}_{Y(\ell)}^{simple}-\widehat{\mathbb{P}}_{\bm{X}}^{(\ell)}G_{Y(\ell)|\bm{X}}\big{)}$
$+\mathrm{Cov}\big{(}\widehat{\mathbb{P}}_{\bm{X}}G_{Y(k)|\bm{X}},\widehat{\mathbb{P}}_{\bm{X}}G_{Y(\ell)|\bm{X}}\big{)},$
where it can be shown that the first and second terms on the right-hand side are equal zero, due to the fact that $\mathbb{E}\big{[}\widehat{\mathbb{F}}_{Y(k)}^{simple}(y)-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}G_{Y(k)|\bm{X}}|\bm{X}_{1},\dots,\bm{X}_{n}\big{]}=0$. Furthermore, under the random sample assumption, we have that $\mathrm{Cov}\big{(}\widehat{\mathbb{P}}_{\bm{X}}G_{Y(k)|\bm{X}},\widehat{\mathbb{P}}_{\bm{X}}G_{Y(\ell)|\bm{X}}\big{)}=n^{-1}\mathrm{Cov}\big{(}G_{Y(k)|\bm{X}},G_{Y(\ell)|\bm{X}}\big{)}$. Thus, we can prove the equality in (A.6).
Next, we compare the variance-covariance matrices of the simple and regression-adjusted estimators. Define $\sigma_{k,\ell}:=\mathrm{Cov}\big{(}G_{Y(k)|\bm{X}},G_{Y(\ell)|\bm{X}}\big{)}$ for $k,\ell\in\mathcal{K}$. By applying the result from part (a) of this theorem and the one in (A.6), we are able to show that
$n\big{\{}\mathrm{Var}\big{(}\widehat{\gamma}^{simple}\big{)}-\mathrm{Var}\big{(}\widetilde{\gamma}\big{)}\big{\}}$ $=\left[\begin{array}[]{cccc}\frac{1-\pi_{1}}{\pi_{1}}\sigma_{11},&-\sigma_{12},&\dots,&-\sigma_{1K}\\ -\sigma_{21},&\frac{1-\pi_{2}}{\pi_{2}}\sigma_{22},&\dots,&-\sigma_{2K}\\ \vdots&\vdots&\ddots&\vdots\\ -\sigma_{K1},&-\sigma_{K2},&\dots,&\frac{1-\pi_{K}}{\pi_{K}}\sigma_{KK}\end{array}\right]+o(1)$
$=\mathbb{E}\Big{[}\big{(}\gamma(\bm{X})-\mathbb{E}[\gamma(\bm{X})]\big{)}A\big{(}\gamma(\bm{X})-\mathbb{E}[\gamma(\bm{X})]\big{)}^{\top}\Big{]}+o(1),$
where
$\gamma(\bm{X})=[G_{Y(1)|\bm{X}},\dots,G_{Y(K)|\bm{X}}]^{\top}\ \ \ \mathrm{and}\ \ \ A:=\left[\begin{array}[]{cccc}{\pi}_{1}^{-1}-1,&-1,&\dots,&-1\\ -1,&{\pi}_{2}^{-1}-1,&\dots,&-1\\ \vdots&\vdots&\ddots&\vdots\\ -1,&-1,&\dots,&{\pi}_{K}^{-1}-1\end{array}\right].$
Lemma A.1 with $\sum_{k=1}^{K}\pi_{k}=1$ shows that, for an arbitrary vector $v:=(v_{1},\dots,v_{K})^{\top}\in\mathbb{R}^{k}$,
$v^{\top}\big{(}\gamma(\bm{X})-\mathbb{E}[\gamma(\bm{X})]\big{)}A\big{(}\gamma(\bm{X})-\mathbb{E}[\gamma(\bm{X})]\big{)}^{\top}v$
$\ \ =\sum_{k\in\mathcal{K}}\frac{v_{k}^{2}\big{(}G_{Y(k)|\bm{X}}-\mathbb{E}[G_{Y(k)|\bm{X}}]\big{)}^{2}}{\pi_{k}}-\bigg{(}\sum_{k\in\mathcal{K}}v_{k}\big{(}G_{Y(k)|\bm{X}}-\mathbb{E}[G_{Y(k)|\bm{X}}]\big{)}\bigg{)}^{2}$
$\ \ \ =\frac{1}{2}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell\in\mathcal{K}\\ \ell\neq k\end{subarray}}\frac{\big{\{}v_{k}\big{(}G_{Y(k)|\bm{X}}-\mathbb{E}[G_{Y(k)|\bm{X}}]\big{)}{\pi}_{\ell}-v_{\ell}\big{(}G_{Y(\ell)|\bm{X}}-\mathbb{E}[G_{Y(\ell)|\bm{X}}]\big{)}{\pi}_{k}\big{\}}^{2}}{{\pi}_{k}{\pi}_{\ell}}.$
It follows that
$v^{\top}\big{\{}\mathrm{Var}\big{(}\widehat{\gamma}^{simple}\big{)}-\mathrm{Var}\big{(}\widetilde{\gamma}\big{)}\big{\}}v=\frac{1}{2}\sum_{k\in\mathcal{K}}\sum_{\begin{subarray}{c}\ell\in\mathcal{K}\\ \ell\neq k\end{subarray}}\frac{\mathrm{Var}\Big{(}v_{k}G_{Y(k)|\bm{X}}{\pi}_{\ell}-v_{\ell}G_{Y(\ell)|\bm{X}}{\pi}_{k}\Big{)}}{{\pi}_{k}{\pi}_{\ell}}+o(n^{-1}).$
Because $\mathrm{Var}\big{(}v_{k}G_{Y(k)|\bm{X}}{\pi}_{\ell}-v_{\ell}G_{Y(\ell)|\bm{X}}{\pi}_{k}\big{)}\geq 0$ for any $k,\ell\in\mathcal{K}$ with $k\neq\ell$, the above equality implies the desired positive semi-definiteness result.
Furthermore, the positive definite result holds when $\mathrm{Var}\big{(}v_{k}G_{Y(k)|\bm{X}}{\pi}_{\ell}-v_{\ell}G_{Y(\ell)|\bm{X}}{\pi}_{k}\big{)}>0$ for any $v\in\mathbb{R}^{K}$ with $v\neq 0$ and for any $k,\ell\in\mathcal{K}$ with $k\neq\ell$. Because $v\in\mathbb{R}^{K}$ is chosen arbitrarily except $v\neq 0$ and $\pi_{k}\in(0,1)$ for all $k\in\mathcal{K}$, the condition for the positive definiteness can be written as $\mathrm{Var}\big{(}G_{Y(k)|\bm{X}}-r\cdot G_{Y(\ell)|\bm{X}}\big{)}>0$ for any $r\in\mathbb{R}$ and for any $k,\ell\in\mathcal{K}$ with $w\neq\ell$. ∎
A.2 Distributional Regression
In this subsection, we present theoretical results of the distributional regression estimator and the conditional distribution estimator. To obtain the asymptotic result in the subsequent section, establishing the convergence rate of the conditional distribution estimator is sufficient. However, for completeness, we also provide their limit distributions. We first obtain a preliminary lemma, which establishes the Donskerness of the first derivative of the log likelihood.
Lemma A.2.
Suppose that Assumptions 1 and 2 hold. Then, the function classes {^$\{\nabla\widehat{\ell}_{k}(\beta_{k};y):(\beta_{k},y)\in\mathcal{B}{\times}\mathcal{Y}\}$ and $\{G(y|\bm{X}):y\in\mathcal{Y}\}$ are Donsker with a square-integrable envelope.
Proof.
We define the function classes, for each $k\in\mathcal{K}$,
$\mathcal{F}_{k}:=\{\bm{x}\mapsto T(\bm{x})^{\top}\beta_{k}:\beta_{k}\in\mathcal{B}\}\ \ \mathrm{and}\ \ \mathcal{G}_{k}:=\{y\mapsto 1\!{\rm l}_{\{y\leq v\}}:v\in\mathcal{Y}\}.$
Lemma 2.6.15 of van der Vaart and Wellner (1996) shows that each of $\{\mathcal{F}_{k}\}_{k=1}^{K}$ and $\{\mathcal{G}_{k}\}_{k=1}^{K}$ are VC-subgraph classes. We can view each element of the regressors $T(\bm{x})\equiv[T^{(1)}(\bm{x}),\dots,T^{(d)}(\bm{x})]^{\top}$ as a function, and then the function classes $\{T(\bm{x})^{(j)}:j=1,\dots,d\}$ are also VC-subgraph classes. Let
$\mathcal{H}:=\cup_{k=1}^{K}\big{\{}[\Lambda(\mathcal{F}_{k})-\mathcal{G}_{k}]R(\mathcal{F}_{k})T^{(j)}(\bm{X})\big{\}}_{j=1}^{p}$
with $\Lambda(\cdot)$ being the link function and $R(\cdot)=\lambda(\cdot)/\{\Lambda(\cdot)[1-\Lambda(\cdot)]\}$. The class $\mathcal{H}$ consists of a Lipschitz transformation of VC-subgraph classes with Lipschitz coefficients and bounded from above by $\|T(\bm{X})\|$ up to some constant factor. Also, an envelop function for $\mathcal{H}$ is bounded from above by $\|T(\bm{X})\|$ up to some constant factor and thus is square integrable under Assumption A1. A Lipschitz composition of a Donsker class is a Donsker class. See 19.20 of van der Vaart (2000). Thus, the desired result for the first class is obtained. A simpler argument prove the desired conclusion for the other class. ∎
The following lemma obtains an asymptotic linear representation for the estimator of parameters in the distributional regression.
Lemma A.3.
Suppose that Assumptions 1 -2 hold. Then, for each $k\in\mathcal{K}$, we have, uniformly in $y\in\mathcal{Y}$,
$\sqrt{n_{k}}\big{(}\widehat{\beta}_{k}(y)-\beta_{k}(y)\big{)}=-H_{k}(y)^{-1}n_{k}^{-1/2}\sum_{i=1}^{n}W_{i,k}\nabla\ell_{i}\big{(}\beta_{1}(y),y\big{)}+o_{p}(1).$
Proof.
Let $k\in\mathcal{K}$ be fixed. To simplify notation, we define a localized objective function,
$\widehat{Q}(\delta;y):=\frac{1}{n_{k}}\sum_{i=1}^{n}W_{i,k}\big{\{}\ell_{i,y}\big{(}\beta_{k}(y)+n_{k}^{-1/2}\delta\big{)}-\ell_{i,y}\big{(}\beta_{k}(y)\big{)}\big{\}}.$
Then, we can write the estimator $\widehat{\delta}(y):=\sqrt{n_{k}}\big{(}\widehat{\beta}_{k}(y)-\beta_{k}(y)\big{)}$ as the solution for $\max_{\delta\in\mathbb{R}^{d}}\widehat{Q}(\delta;y)$.
Let $M$ be a finite positive constant. Under Assumption A2, the objective function $\delta\mapsto\widehat{Q}(\delta;y)$ is twice continuously differentiable, we can show that, for each fixed $\delta$ with $\|\delta\|\leq M$, uniformly in $y\in\mathcal{Y}$,
$n_{k}\widehat{Q}(\delta;y)=\delta^{\top}\sqrt{n_{k}}\nabla\widehat{Q}(0;y)+\frac{1}{2}\delta^{\top}\nabla^{2}\widehat{Q}(0;y)\delta+o(n_{k}^{-1}\|\delta\|^{2}).$
Also, an application of the uniform law of large numbers for the Hessian matrix shows that $\nabla^{2}\widehat{Q}(0;y)\to^{p}H_{k}(y)$ uniformly in $y\in\mathcal{Y}$. Thus, for each $\delta$ with $\|\delta\|\leq M$, we can show that $\sup_{y\in\mathcal{Y}}\big{|}\widehat{Q}(\delta;y)-\widetilde{Q}(\delta;y)\big{|}=o_{p}(n^{-1})$, where
$n_{k}\widetilde{Q}(\delta;y)=\delta^{\top}\sqrt{n_{k}}\nabla\widehat{Q}(0;y)+\frac{1}{2}\delta^{\top}H_{k}(y)\delta.$
The convexity lemma (see Pollard, 1991; Kato, 2009) extends the point-wise convergence with respect to $\delta$ to the uniform converges and thus, under Assumption A2,
$\sup_{y\in\mathcal{Y}}\sup_{\delta:\|\delta\|\leq M}\big{|}\widehat{Q}(\delta;y)-\widetilde{Q}(\delta;y)\big{|}=o_{p}(n_{k}^{-1}).$ (A.7)
Because $\widetilde{Q}(\delta;y)$ is maximized at $\tilde{\delta}(y):=-H_{k}(y)^{-1}\sqrt{n_{k}}\nabla\widehat{Q}(0;y)$, simple algebra can show that, for any $\delta$ and for some constant $c>0$,
$\widetilde{Q}\big{(}\tilde{\delta}(y);y\big{)}-\widetilde{Q}\big{(}\delta;y\big{)}$ $=-\frac{1}{2n_{k}}\big{(}\tilde{\delta}(y)-\delta\big{)}^{\top}H_{k}(y)\big{(}\tilde{\delta}(y)-\delta\big{)}$
$\geq\frac{c}{2n_{k}}\|\tilde{\delta}(y)-\delta\|^{2},$ (A.8)
where the last inequality is due to that $H_{k}(y)$ is negative definite under Assumption A4. For any subset $D$ including $\tilde{\delta}(y)$, an application of the triangle inequality obtains
$2\sup_{\delta\in D}|\widehat{Q}(\delta;y)-\widetilde{Q}(\delta;y)|$ $\geq\sup_{\delta\in D}\big{\{}\widetilde{Q}\big{(}\tilde{\delta}(y);y\big{)}-\widetilde{Q}\big{(}\delta;y\big{)}\big{\}}$
$\ \ \ -\sup_{\delta\in D}\big{\{}\widehat{Q}\big{(}\tilde{\delta}(y);y\big{)}-\widehat{Q}(\delta;y)\}.$ (A.9)
Let $\eta>0$ be an arbitrary constant. Because of the concavity under Assumption A2, difference quotients satisfy that, for any $\lambda>\eta$ and for any $v\in S^{k}$ with the unit sphere $S^{k}\subseteq\mathbb{R}^{k}$,
$\frac{\widehat{Q}\big{(}\tilde{\delta}(y)+\eta v;y\big{)}-\widehat{Q}\big{(}\tilde{\delta}(y);y\big{)}}{\eta}\geq\frac{\widehat{Q}\big{(}\tilde{\delta}(y)+\lambda v;y\big{)}-\widehat{Q}\big{(}\tilde{\delta}(y);y\big{)}}{\lambda}.$
This inequality with a set $D_{n,\eta}(y):=\{\delta\in\mathbb{R}^{k}:\|\delta-\tilde{\delta}(y)\|\leq\eta\}$ implies that, given the event $\{\sup_{y\in\mathcal{Y}}\|\hat{\delta}(y)-\tilde{\delta}(y)\|\geq\eta\}$, we have, for any $y\in\mathcal{Y}$,
$\sup_{\delta\in D_{n_{k},\eta}(y)}\widehat{Q}(\delta;y)-\widehat{Q}\big{(}\tilde{\delta}(y);y\big{)}\geq 0,$ (A.10)
where the last inequality is due to that $\widehat{Q}\big{(}\hat{\delta}(y);y\big{)}-\widehat{Q}\big{(}\tilde{\delta}(y);y\big{)}\geq 0$, by definition of $\hat{\delta}(y)$. It follows from (A.8)-(A.10) that, given the event $\{\sup_{y\in\mathcal{Y}}\|\hat{\delta}(y)-\tilde{\delta}(y)\|\geq\eta\}$, we have
$\sup_{\delta\in D_{n_{k},\eta}(y)}|\widehat{Q}(\delta;y)-\widetilde{Q}(\delta;y)|\geq\frac{c}{4n_{k}}\eta^{2}.$
Lemma A.2 shows that $\nabla\widehat{Q}(0;y)=\nabla\widehat{\ell}_{k}\big{(}\beta_{k}(y);y\big{)}$ is Donsker. Applying the Donsker property for $\tilde{\delta}(y):=-H_{k}(y)^{-1}\sqrt{n_{k}}\nabla\widehat{Q}(0;y)$, we can show that, for any $\xi>0$, there exists a constant $C$ such that $\Pr(\sup_{y\in\mathcal{Y}}\|\tilde{\delta}(y)\|\geq C)\leq\xi$ for sufficiently large $n_{k}$. Thus, the above display implies that
$\Pr\bigg{(}\sup_{y\in\mathcal{Y}}\|\hat{\delta}(y)-\tilde{\delta}(y)\|\geq\eta\bigg{)}\leq\Pr\bigg{(}\sup_{y\in\mathcal{Y}}\sup_{\delta:\|\delta\|\leq\eta+C}\big{|}\widehat{Q}(\delta)-\widetilde{Q}(\delta)\big{|}>\frac{c}{4n_{k}}\eta^{2}\bigg{)}+\xi,$
for sufficiently large $n_{k}$. It follows from (A.7) that the first term on the right side of the above equation converges to 0 as $n_{k}\to\infty$. Thus, we obtain the desired conclusion. ∎
The theorem below present the asymptotic distribution of the estimator of the parameters of the DR model.
Proposition A.1.
Under Assumptions 1 and 2, we have, for each $k\in\mathcal{K}$,
$\sqrt{n_{k}}\big{(}\widehat{\beta}_{k}(\cdot)-\beta_{k}(\cdot)\big{)}\rightsquigarrow\mathbb{V}_{k}(\cdot)\ \ \mathrm{in}\ \ \ell^{\infty}(\mathcal{Y})^{p},$
where $\mathbb{V}_{k}(y)$ is a mean-zero Gaussian process over $\mathcal{Y}$, and its covariance function is given by $H_{k}(y)^{-1}\Sigma_{k}(y,y^{\prime})H_{k}(y^{\prime})^{-1}$ for $y,y^{\prime}\in\mathcal{Y}$ with $\Sigma_{k}(y,y^{\prime}):=\mathbb{E}[W_{i,k}\nabla_{\beta}\ell_{i}\big{(}\beta_{k}(y),y\big{)}\nabla_{\beta}\ell_{i}\big{(}\beta_{k}(y),y\big{)}^{\top}]$.
Proof.
Fix $k\in\mathcal{K}$. In this proof, we consider the case where outcome variable is a continuous random variable, while a simpler proof can apply for the case of discrete outcome variables. Lemma A.3 implies that, uniformly in $y\in\mathcal{Y}$,
$\sqrt{n_{k}}\big{(}\hat{\beta}_{k}(y)-\beta_{k}(y)\big{)}=-H_{k}(y)^{-1}\sqrt{n_{k}}\nabla\widehat{\ell}_{k}\big{(}\beta_{k}(y),y\big{)}+o_{p}(1).$
By the implicit function theorem, we can show that $y\mapsto\beta_{k}(y)$ is differentiable uniformly over $y\in\mathcal{Y}$. Thus, the empirical process ^$\nabla\widehat{\ell}_{k}\big{(}\beta_{k}(y),y\big{)}$ is stochastically equicontinuous over $\mathcal{Y}$. Given iid observations under Assumption A1, the finite dimensional convergence follows from a multivariate central limit theorem. This with the stochastic equicontinuity implies that ^$\nabla\widehat{\ell}_{k}\big{(}\beta_{k}(y),y\big{)}\rightsquigarrow\mathbb{G}_{k}(\cdot)$ in $\ell^{\infty}(\mathcal{Y})^{k}$, where $\mathbb{G}(\cdot)$ is a zero-mean Gaussian process with covariance function $\Sigma_{k}(\cdot,\cdot)$. The desired result follows, after taking $\mathbb{V}_{k}(\cdot)=-H_{k}(\cdot)^{-1}\mathbb{G}_{k}(\cdot)$. ∎
The distribution function estimator, $\widehat{G}_{Y(k)|\bm{X}}$ is a transformation of the estimator $\hat{\beta}_{k}(\cdot)$. More precisely, let $\mathbb{D}:=\ell^{\infty}(\mathcal{Y})^{d}$ and define the map $\phi:\mathbb{D}_{\phi}\subset\mathbb{D}\mapsto\mathbb{S}_{\phi}$, given by
$\phi(\beta)(\bm{x},y):=\Lambda\big{(}T(\bm{x})^{\top}\beta(y)\big{)}.$
Then, we can write
$\widehat{G}_{Y(k)|\bm{X}}=\phi(\widehat{\beta}_{k}).$
It can be shown that the map $\phi$ is Hadamard differentiable at $\beta_{k}\in\mathbb{D}_{\phi}$ tangentially to $\mathbb{D}$ with the Hadamard derivative $b_{k}\mapsto\phi_{\beta_{k}(\cdot)}^{\prime}(b_{k})$, given by
$\phi_{\beta_{k}(\cdot)}^{\prime}(b_{k})(\bm{x},y):=\lambda\big{(}T(\bm{x})^{\top}\beta_{k}(y)\big{)}T(\bm{x})^{\top}b_{k}(y).$
We can obtain the joint asymptotic distribution of the distribution function estimators, applying the functional delta method with the Hadamard derivative in the above display. The result is formally stated in the theorem below.
Proposition A.2.
Under Assumptions 1 and 2, we have,
$\sqrt{n}\left(\begin{array}[]{c}\widehat{G}_{Y(1)|\bm{X}}-G_{Y(1)|\bm{X}}\\ \vdots\\ \widehat{G}_{Y(K)|\bm{X}}-G_{Y(K)|\bm{X}}\end{array}\right)\rightsquigarrow\left[\begin{array}[]{c}\pi_{1}^{-1/2}\phi_{\beta_{1}(\cdot)}^{\prime}\big{(}\mathbb{V}_{1}\big{)}\\ \vdots\\ \pi_{K}^{-1/2}\phi_{\beta_{K}(\cdot)}^{\prime}\big{(}\mathbb{V}_{K}\big{)}\\ \end{array}\right]\ \ \mathrm{in}\ \ell^{\infty}(\mathcal{X}{\times}\mathcal{Y})^{2},$
where $\{\mathbb{V}_{k}\}_{k=1}^{K}$ are the mean-zero Gaussian process defined in Theorem A.1.
Proof.
Consider the map $\phi:\mathbb{D}_{\phi}\subset\mathbb{D}\mapsto\mathbb{S}_{\phi}$, where $\beta\mapsto\phi(\beta)$, given by $\phi(\beta)(\bm{x},y)=\Lambda\big{(}T(\bm{x})^{\top}\beta(y)\big{)}$. Then, we can write $\widehat{G}_{Y(k)|\bm{X}}=\phi\big{(}\hat{\beta}_{k}(\cdot)\big{)}$ and $G_{Y(k)|\bm{X}}=\phi\big{(}\beta_{k}(\cdot)\big{)}$ for $k\in\mathcal{K}$. Under Assumption 2, the map $\phi(\cdot)$ is shown to be Hadamard differentiable at $\beta_{k}(\cdot)$ tangentially to $\mathbb{D}$ with the derivative map $b\mapsto\phi_{\beta_{k}(\cdot)}^{\prime}(b)$ for every $k\in\mathcal{K}$, given by
$\phi_{\beta_{k}(\cdot)}^{\prime}(b)(\bm{x},y)=\lambda\big{(}T(\bm{x})^{\top}\beta_{k}(y)\big{)}T(\bm{x})^{\top}b(y).$
Applying the functional delta method with the result in Theorem A.1, we can show that, for every $k\in\mathcal{K}$,
$\sqrt{n_{k}}\big{(}\widehat{G}_{Y(k)|\bm{X}}-G_{Y(k)|\bm{X}}\big{)}\rightsquigarrow\phi_{\beta_{k}(\cdot)}^{\prime}(\mathbb{B}_{k})\ \ \mathrm{in}\ \ \ell^{\infty}(\mathcal{X}{\times}\mathcal{Y}).$
The Gaussian processes $\{\mathbb{B}_{k}(\cdot)\}_{k=1}^{K}$ are mutually independent under the random sample assumption. Also, we have $n_{k}/n=\pi_{k}+o(1)$. Thus, the desired conclusion follows. ∎
A.3 Proof of Theorem 2
For treatment $k\in\mathcal{K}$, define the empirical processes
$\widehat{\nu}_{\bm{X}}^{(k)}:=\sqrt{n_{k}}\big{(}\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-P_{\bm{X}}^{(k)}\big{)}\ \ \ \mathrm{and}\ \ \ \widehat{\nu}_{Y(k)}^{simple}:=\sqrt{n_{k}}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}-F_{Y(k)}\big{)},$
In the below lemma, we derive the joint limit of the empirical process.
Lemma A.4.
Suppose that Assumption 1 hold. Then, we have,
$\big{\{}\big{(}\widehat{\nu}_{\bm{X}}^{(k)},\widehat{\nu}_{Y(k)}^{simple}\big{)}\big{\}}_{k=1}^{K}\rightsquigarrow\big{\{}\big{(}\mathbb{B}_{\bm{X}}^{(k)},\mathbb{B}_{Y(k)}\big{)}\big{\}}_{k=1}^{K}\ \ \ \mathrm{in}\ \ \ell^{\infty}(\mathcal{X}\times\mathcal{Y})^{K}.$
Here, $\mathbb{B}_{\bm{X}}^{(k)}$ and $\mathbb{B}_{Y(k)}^{(k)}$ are mean-zero Gaussian processes for $k\in\mathcal{K}$. Given an arbitrary function $m(y,x)\in\ell^{\infty}(\mathcal{Y}{\times}\mathcal{X})$ having the Donsker property, the covariance function of the pair $\big{(}\mathbb{B}_{\bm{X}}^{(k)}m(\cdot,\bm{X}),\mathbb{B}_{Y(k)},\big{)}$ is given by, for any $y,y^{\prime}\in\mathcal{Y}$,
$\mathrm{Cov}\Big{(}\big{(}1\!{\rm l}_{\{Y(k)\leq y\}},m(y,\bm{X})\big{)},\big{(}1\!{\rm l}_{\{Y(k)\leq y^{\prime}\}},m(y^{\prime},\bm{X})\big{)}\Big{)}.$
Moreover, the pairs $\{(\mathbb{B}_{\bm{X}}^{(k)},\mathbb{B}_{Y(k)})\}_{k=1}^{K}$ are statistically independent across treatments.
Proof.
Fix $k\in\mathcal{K}$. By definition, we can write $\widehat{\nu}_{Y(k)}^{simple}=\sqrt{n_{k}}\sum_{i=1}^{n}W_{ik}\cdot\big{(}1\!{\rm l}_{\{Y(k)\leq y\}}-F_{Y(k)}(y)\big{)}$ and it can be shown that the class $\{1\!{\rm l}_{\{Y(k)\leq y\}}:y\in\mathcal{Y}\}$ is Donsker. The Donskerness of $\{1\!{\rm l}_{\{Y(k)\leq y\}}:y\in\mathcal{Y}\}$ and $\{m(y,X):y\in\mathcal{Y}\}$ implies the stochastic equicontinuity. Thus, it remains to show the finite dimensional convergence, which follows from a simple application of the multivariate central limit theorem, given the iid observations under Assumption 1.
Under Assumption 1, the observations of the treatment groups are mutually independent, thereby yielding the pairwise independence of the limit processes. ∎
Proof of Theorem 2.
For treatment $k\in\mathcal{K}$, we have that $\widehat{\mathbb{F}}_{Y(k)}^{simple}=\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\widehat{G}_{Y(k)|\bm{X}}$, provided that the link function is of the canonical class, and we can write
$\widehat{\mathbb{F}}_{Y(k)}=\widehat{\mathbb{F}}_{Y(k)}^{simple}+\big{(}\widehat{\mathbb{P}}_{\bm{X}}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\big{)}\widehat{G}_{Y(k)|\bm{X}}.$
This yields
$\sqrt{n}\big{(}\widehat{\mathbb{F}}_{Y(k)}-F_{Y(k)}\big{)}=\frac{\sqrt{n}}{\sqrt{n_{k}}}\sqrt{n_{k}}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}-F_{Y(k)}\big{)}+\sqrt{n}\big{(}\widehat{\mathbb{P}}_{\bm{X}}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\big{)}\widehat{G}_{Y(k)|\bm{X}}.$
The pre-treatment variable $\bm{X}$ shares the population distribution $P_{\bm{X}}$ across treatment groups and also ^$\widehat{\mathbb{P}}_{\bm{X}}=\sum_{\ell=1}^{K}(n_{\ell}/n)\widehat{\mathbb{P}}_{\bm{X}}^{(\ell)}$. Thus, we can write
^$\widehat{\mathbb{P}}_{\bm{X}}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}$ $=\sum_{\ell=1}^{K}\frac{n_{\ell}}{n}(\widehat{\mathbb{P}}_{\bm{X}}^{(\ell)}-P_{\bm{X}})-(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-P_{\bm{X}})$
$=\frac{1}{\sqrt{n}}\bigg{(}\sum_{\ell=1}^{K}\frac{\sqrt{n_{\ell}}}{\sqrt{n}}\widehat{\nu}_{\bm{X}}^{(\ell)}-\frac{\sqrt{n}}{\sqrt{n_{k}}}\widehat{\nu}_{\bm{X}}^{(k)}\bigg{)}$
The empirical processes $\{\widehat{\nu}_{\bm{X}}^{(\ell)}(\cdot)\}_{\ell\in\mathcal{K}}$ are stochastically equicontinuous. Thus, it can be shown that $\widehat{\nu}_{\bm{X}}^{(\ell)}\big{(}\widehat{G}_{Y(k)|\bm{X}}-G_{Y(k)|\bm{X}}\big{)}=o_{p}(1)$ for every $\ell\in\mathcal{K}$, given the result that $\sqrt{n}\big{(}\widehat{G}_{Y(k)|\bm{X}}-G_{Y(k)|\bm{X}}\big{)}=O_{p}(1)$ by Theorem A.2. Thus, we have
$\sqrt{n}(\widehat{\mathbb{P}}_{\bm{X}}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)})\widehat{G}_{Y(k)|\bm{X}}=\bigg{(}\sum_{\ell=1}^{K}\sqrt{\pi_{\ell}}\widehat{\nu}_{\bm{X}}^{(\ell)}-\frac{1}{\sqrt{\pi_{k}}}\widehat{\nu}_{\bm{X}}^{(k)}\bigg{)}G_{Y(k)|\bm{X}}+o_{p}(1),$
which leads to
$\sqrt{n}\big{(}\widehat{\mathbb{F}}_{Y(k)}-F_{Y(k)}\big{)}=\frac{1}{\sqrt{\pi_{k}}}\widehat{\nu}_{Y(k)}^{simple}+\bigg{(}\sum_{\ell=1}^{K}\sqrt{\pi_{\ell}}\widehat{\nu}_{\bm{X}}^{(\ell)}-\frac{1}{\sqrt{\pi_{k}}}\widehat{\nu}_{\bm{X}}^{(k)}\bigg{)}G_{Y(k)|\bm{X}}+o_{p}(1).$
It follows from Lemma A.2 that
$\sqrt{n}\left(\begin{array}[]{c}\widehat{F}_{Y(1)}-F_{Y(1)}\\ \vdots\\ \widehat{F}_{Y(K)}-F_{Y(K)}\end{array}\right)\rightsquigarrow\mathbb{B}:=\left[\begin{array}[]{c}\mathbb{B}_{1}\\ \vdots\\ \mathbb{B}_{K}\\ \end{array}\right]\ \ \mathrm{in}\ \ell^{\infty}(\mathcal{X}{\times}\mathcal{Y})^{2},$
with
$\mathbb{B}_{k}:=$ $\frac{1}{\sqrt{\pi_{k}}}\mathbb{B}_{Y(k)}+\bigg{(}\sum_{\ell=1}^{K}\sqrt{\pi_{\ell}}\mathbb{B}_{\bm{X}}^{(\ell)}-\frac{1}{\sqrt{\pi_{k}}}\mathbb{B}_{\bm{X}}^{(k)}\bigg{)}G_{Y(k)|\bm{X}}.$
As $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ and $\Psi^{PTE}_{k,k^{\prime},h}(\cdot)$ are linear transformations, the conclusion follows. ∎
A.4 Proof of Theorem 3
For the validity of the resampling method, we introduce the notion of conditional weak convergence in probability, following van der Vaart and Wellner (1996).
Definition 1.
[…]
where $\mathbb{E}_{n}$ is the expectation with respect to $\{S_{i}\}_{i=1}^{n}$ conditional on the original sample.
Using the bootstrap observations, we define the empirical processes, for treatment $k\in\mathcal{K}$,
$\widehat{\nu}_{\bm{X}}^{(k)\ast}:=\sqrt{n_{k}}\big{(}\widehat{\mathbb{P}}_{\bm{X}}^{(k)\ast}-\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\big{)}\ \ \mathrm{and}\ \ \widehat{\nu}_{Y(k)}^{simple\ast}:=\sqrt{n_{k}}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple\ast}-\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)},$
where ^$\widehat{\mathbb{P}}_{\bm{X}}^{(k)\ast}:=n_{k}^{-1}\sum_{i=1}^{n}S_{i}\cdot W_{i,k}\cdot\delta_{\bm{X}_{i}}.$
Proof of Theorem 3.
Applying a similar argument in the proof of Theorem 2, we can show that
$\sqrt{n}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{\ast}-\widehat{\mathbb{F}}_{Y(k)}\big{)}=\frac{1}{\sqrt{\pi_{k}}}\widehat{\nu}_{Y(k)}^{simple\ast}+\bigg{(}\sum_{\ell=1}^{K}\sqrt{\pi_{\ell}}\widehat{\nu}_{\bm{X}}^{(\ell)\ast}-\frac{1}{\sqrt{\pi}}\widehat{\nu}_{\bm{X}}^{(k)\ast}\bigg{)}G_{Y(k)|\bm{X}}+o_{p}(1),$
in $\ell^{\infty}(\mathcal{Y})$, for every $k\in\mathcal{K}$. Because Lemma A.2 shows that the function classes considered in the above equation are Donsker, Theorem 3.6.13 of van der Vaart and Wellner (1996) with the functional delta method yields
[…]
in $\ell^{\infty}(\mathcal{Y})$. Furthermore, the transformations $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ and $\Psi^{PTE}_{k,k^{\prime},h}(\cdot)$ are linear in the regression-adjusted estimator, the desired conclusion follows. ∎
Appendix B Additional Simulation Results
The remaining simulation results are presented in Figures 5-10. The key insights discussed in Section 4 remain consistent across these figures.
Figure 5: Performance metrics of simple and regression-adjusted DTE estimators
Note: (DGP1, continuous outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
Figure 6: Performance metrics of simple and regression-adjusted DTE estimators
Note: (DGP2, continuous outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
Figure 7: Performance metrics of simple and regression-adjusted DTE estimators
Note: (DGP2, continuous outcome, $\pi_{1}=0.5$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
Figure 8: Performance metrics of simple and regression-adjusted DTE estimators
Note: (DGP3, discrete outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
Figure 9: Performance metrics of simple and regression-adjusted DTE estimators
Note: (DGP4, discrete outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
Figure 10: Performance metrics of simple and regression-adjusted DTE estimators
Note: (DGP4, discrete outcome, $\pi_{1}=0.5$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.